Eiwoh2:

Suppose C and D represent two different school populations where C > D and C and D must be greater than 0. Which of the following expressions is the largest? Explain why. Show all work necessary. (C + D)^2 2(C + D) C2 + D^2 C2 − D^2

Eiwoh2:

@Shadow @dude I think the largest the third one, but I don't know how to prove it.

mhchen:

Well you can compare them. (C + D)^2 = C^2 + 2CD + D^2 2(C + D) = 2C + 2D Did you teacher tell you that C^2 is bigger than 2C?

Eiwoh2:

Well, i kinda figured that out with the fact that C^2 would be more equal than just C. Plus, the equation has a plus sign.

mhchen:

Okay so like C^2 = C * C 2C = 2 * C So if C is bigger than 2 then C^2 is bigger than 2C otherwise it's equal. So we'll just say C^2 can be greater than 2C Now it says C is bigger than D. So C^2 is bigger than D^2. Out of all of them. Only the first equation has C^2 and D^2 while the rest only has D^2 or no exponents at all. Understand?

Eiwoh2:

OH. So the first one would be the biggest, since it combines C and D and squares it...ok. I get it like that. XD